OpenStudy (anonymous):

John is makinga shake by mixing two different protein powders, measured in ounces. The strawberry-flavored powder has 4 grams of protein per ounce. The banana-flavored powder has 3 grams of protein per ounce. He wants the drink to have a total of 6 ounces of powder and contain 22 grams of protein.

OpenStudy (johnweldon1993):

Alright so the total OUNCES he wants from a mixture of both...is 6 Let strawberry = s and banana = b s + b = 6 We also know there are 4 grams per ounce in the strawberry and 3 grams per ounce of the banana...and he wants a total of 22 grams of protein...so 4s + 3b = 22 We have a system of equations here s + b = 6 4s + 3b = 22

OpenStudy (johnweldon1993):

Do you know how to solve these types of problems?

OpenStudy (anonymous):

yes you just replace the s with 3 and the b with 3, right?

OpenStudy (johnweldon1993):

Not quite

OpenStudy (johnweldon1993):

We want to solve that top equation for 's'

OpenStudy (anonymous):

ohhh..

OpenStudy (johnweldon1993):

so s + b = 6 s = 6 - b Now we plug that in for 's' in our second equation 4s + 3b = 22 becomes 4(6 - b) + 3b = 22 24 - 4b + 3b = 22 -b = -2 b = 2 So this means he wants 2 ounces of banana And if he wants 2 banana...he wants s + b = 6 s + 2 = 6 s = 4 he would want 4 ounces of strawberry... And to check...plug b = 2 and s = 4 into equation 2 4s + 3b = 22 4(4) + 3(2) = 22 16 + 6 = 22 22 = 22 \(\large \checkmark\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you get -b

OpenStudy (johnweldon1993):

s + b = 6 We wanted to solve this...for 's' ....because we wanted to be able to plug in something for 's' in the second equation.. so I subtracted 'b' from both sides...this made 's' alone on the left side and gave us s = 6 - b

OpenStudy (anonymous):

ok thank you

OpenStudy (johnweldon1993):

No problem!